# Manifold Learning by Mixture Models of VAEs for Inverse Problems

### v2

@@ 4.2 Retractions for Learned Charts
In order to apply this gradient scheme for a learned manifold given by the learned mappings $(\mathcal{D}_{k},\mathcal{E}_{k})_{k=1}^{K}$, we consider two types of retractions. We introduce them and show that they are indeed retractions in the following lemmas. The first one generalizes the idea from [2, Lemma 4, Proposition 5] of moving along the tangent vector in $\mathbb{R}^{n}$ and reprojecting onto the manifold. However, the results from [2] are based on the orthogonal projection, which is hard or even impossible to compute. Thus, we replace it by some more general projection $\pi$. In our applications, $\pi$ will be chosen as in Remark 3, i.e., we set $\pi(x)=\mathcal{D}_{k}(\mathcal{E}_{k}(x)).$
Lemma 7.
@@
Proof.
The property $R_{x}(0)=x$ is directly clear from the definition of $R_{x}$. Now let $h\in T_{x}\mathcal{M}\subseteq\mathbb{R}^{n}$ and $\gamma_{h}\colon(-\epsilon,\epsilon)\to\mathcal{M}$ be a differentiable curve with $\gamma_{h}(0)=x$ and $\dot{\gamma}_{h}(0)=h$. As ${\pi}_{|U}$ is the identity on $\mathcal{M}$, we have by the chain rule that
@@
The second retraction uses the idea of changing to local coordinates, moving into the gradient direction by using the local coordinates and then going back to the manifold representation. Note that similar constructions are considered in [1, Section 4.1.3]. However, as we did not find an explicit proof for the lemma, we give it below for the sake of completeness.
Lemma 8.
@@
Proof.
The property $R_{x}(0)=x$ is directly clear from the definition of $R_{x}$. Now let $h\in T_{0}(T_{x}\mathcal{M})=T_{x}\mathcal{M}\subseteq\mathbb{R}^{n}$. By (15), we have that there exists some $y\in\mathbb{R}^{d}$ such that $h=Jy$. Then, we have by the chain rule that
@@ 4.3 Gradient Descent on Learned Manifolds
By Lemma 7 and 8, we obtain that the mappings
@@
Remark 9 (Descent algorithm).
By construction, Algorithm 3 is a descent algorithm. That is, for a sequence $(x_{n})_{n}$ generated by the algorithm it holds that $F(x_{n+1})\leq F(x_{n})$. With the additional assumption that $F$ is bounded from below, we have that $(F(x_{n}))_{n}$ is a bounded descending and hence convergent sequence. However, this does neither imply convergence of the iterates $(x_{n})_{n}$ themselves nor optimality of the limit of $(F(x_{n}))_{n}$. For more details on the convergence of line-search algorithms on manifolds we refer to [1, Section 4].
@@ 5 Numerical Examples
Remark 10 (Dimension of the Manifold).
For all our numerical experiments, we assume that the dimension $d$ of the data manifold is known. This assumption might be violated for practical applications. However, there exist several methods in the literature to estimate the dimension of a manifold from data, see e.g., [14, 23, 37, 68]. We are aware that dimension estimation of high dimensional datasets is a hard problem which cannot be considered as completely solved so far. In particular, most of these algorithms make assumptions on the distribution of the given data points. Even though it is an active area of research, it is not the scope of this paper to test, benchmark or develop such algorithms. Similarly, combining them with our mixture of VAEs is left for future research.
@@ 6 Mixture of VAEs for Inverse Problems
In many inverse problems, the unknown $x$ can be modeled as an element of a low-dimensional manifold $\mathcal{M}$ in $\mathbb{R}^{n}$ ([4, 10, 20, 55, 77, 87, 74, 3]), and this manifold can be represented by the mixture of VAEs as explained in Section 3. Thus, the solution of (18) can be found by optimizing the function
@@
Throughout these experiments we consider images of size $128\times 128$ and use the architecture from Section 3.2 with $L=3$. Starting with the latent dimension $d$, the mapping $A_{1}\colon\mathbb{R}^{d}\to\mathbb{R}^{32^{2}}$ fills up the input vector $x$ with zeros up to the size $32^{2}$, i.e., we set $A_{1}(x)=(x,0)$. The invertible neural network $T_{1}\colon\mathbb{R}^{32^{2}}\to\mathbb{R}^{32^{2}}$ consists of $3$ invertible blocks, where the subnetworks $s_{i}$ and $t_{i}$, $i=1,2$ are dense feed-forward networks with two hidden layers and $64$ neurons. Afterwards, we reorder the dimensions to obtain an image of size $32\times 32$. The mappings $A_{2}\colon\mathbb{R}^{32\times 32}\to\mathbb{R}^{32\times 32\times 4}$ and $A_{3}\colon\mathbb{R}^{64\times 64}\to\mathbb{R}^{64\times 64\times 4}$ copy each channel $4$ times. Then, the generalized inverses $A_{2}^{\dagger}\colon\mathbb{R}^{32\times 32\times 4}\to\mathbb{R}^{32\times 32}$ and $A_{3}^{\dagger}\colon\mathbb{R}^{64\times 64\times 4}\to\mathbb{R}^{64\times 64}$ from (11) are given by taking the mean of the four channels of the input. The invertible neural networks $T_{2}\colon\mathbb{R}^{32\times 32\times 4}\to\mathbb{R}^{64\times 64}$ and $T_{3}\colon\mathbb{R}^{64\times 64\times 4}\to\mathbb{R}^{128\times 128}$ consist of $3$ invertible blocks, where the subnetworks $s_{i}$ and $t_{i}$, $i=1,2$ are convolutional neural networks with one hidden layer and $64$ channels. After these three coupling blocks we use an invertible upsampling ([35]) to obtain the correct output dimension. For the normalizing flow in the latent space, we use an invertible neural network with three blocks, where the subnetworks $s_{i}$ and $t_{i}$, $i=1,2$ are dense feed-forward networks with two hidden layers and $64$ neurons.
@@ 6.2 Electrical Impedance Tomography
Finally, we consider the highly non-linear and ill-posed inverse problem of electrical impedance tomography (EIT, [28]), which is also known in the mathematical literature as the Calderón problem ([12, 39, 76]). EIT is a non-invasive, radiation-free method to measure the conductivity of a tissue through electrodes placed on the surface of the body. More precisely, electrical currents patterns are imposed on some of these electrodes and the resulting voltage differences are measured on the remaining ones. Although harmless, the use of this modality in practice is very limited because the standard reconstruction methods provide images with very low spatial resolution. This is an immediate consequence of the severe ill-posedness of the inverse problem ([5, 73]).
Classical methods for solving this inverse problem include variational-type methods ([29]), the Lagrangian method ([27]), the factorization method ([22, 63]), the D-bar method ([89]), the enclosure method ([56]), and the monotonicity method ([93]). Similarly as many other inverse problems, deep learning methods have had a big impact on EIT. For example, the authors of [38] propose an end-to-end neural network that learns the forward map $\mathcal{G}$ and its inverse. Moreover, deep learning approaches can be combined with classical methods, e.g, by post processing methods ([48, 47]) or by variational learning algorithms ([87]).
@@
Remark 11 (Dimension and topology of the data manifold).
@@
A slightly more general version of this manifold was considered in [3], where Lipschitz stability is proven for a related inverse boundary value problem restricted to the manifold. Other types of inclusions (with unknown locations), notably polygonal and polyhedral inclusions, have been considered in the literature ([18, 17, 11]). The case of small inclusions is discussed in [7].
@@
Since the PDE (20) is linear, the map $\mathcal{G}(\gamma)$ is linear. However, the forward map $\gamma\mapsto\mathcal{G}(\gamma)$ is nonlinear in $\gamma$, and so is the corresponding inverse problem. The map $\mathcal{G}$ is continuously differentiable, and its Fréchet derivative (see [49]) is given by $[\nabla\mathcal{G}(\gamma)](\sigma)(g)=w_{g}\big{|}_{\partial\Omega}$, where $w_{g}\in H^{1}(\Omega)$ is the unique weak solution with zero boundary mean of
@@
In our implementations, we discretize the linear mappings $G(\gamma)$ by restricting them to a finite dimensional subspace spanned by a-priori fixed boundary functions $g_{1},\dots,g_{N}\in H^{-\frac{1}{2}}_{\diamond}(\partial\Omega)$. Then, following [19, eqt. (2.2)], we reconstruct the conductivity by minimizing the semi-discrete functional
@@
where $(u_{\text{true}})_{g_{n}}$ is the observed data. In our discrete setting, we represent the conductivitiy $\gamma$ by a piecewise constant function $\gamma=\sum_{m=1}^{M}\gamma_{m}\mathbbm{1}_{T_{m}}$ on a triangulation $(T_{m})_{m=1,\dots,M}$. Then, following [19, eqt. (2.20)], the derivative of (21) with respect to $\gamma$ is given by
@@
In our experiments the domain $\Omega$ is given by the unit square $[0,1]^{2}$. For solving the PDEs (20) and (23), we use a finite element solver from the DOLFIN library ([70]). We employ meshes that are coarser in the middle of $\Omega$ and finer close to the boundary. To simulate the approximation error of the meshes, and to avoid inverse crimes, we use a fine mesh to generate the observation and a coarser one for the reconstructions. We use $N=15$ boundary functions, which are chosen as follows. We divide each of the four edges of the unit square $[0,1]^{2}$ into 4 segments and denote by $b_{1},\dots,b_{16}$ the functions that are equal to $1$ on one of these segments and 00 otherwise. Then, we define the boundary functions as $g_{n}=\sum_{i=1}^{16}a_{n,i}b_{i}$, where the matrix $A=(a_{n,i})_{n=1,\dots,15,i=1,\dots,16}$ is the $16\times 16$ Haar matrix without the first row. More precisely, the rows of $A$ are given by the rows of the matrices $2^{-k/2}(\mathrm{Id}_{2^{4-k}}\otimes(1,-1)\otimes e_{2^{k-1}}^{\mathrm{T}})$ for $k=1,\dots,4$, where $\otimes$ is the Kronecker product and $e_{j}\in\mathbb{R}^{j}$ is the vector where all entries are $1$.
@@ Results.
The reconstructions for 20 different ground truths are visualized in Figure 9(d). We observe that both models capture the ground truth structure in most cases, but also fail sometimes. Nevertheless, the reconstructions with the mixture of two VAEs recover the correct structure more often and more accurately than the single VAE, which can be explained by the better coverage of the data manifold. To quantify the difference more in detail, we rerun the experiment with $200$ different ground truth images and compare the results with one and two generators using the PSNR and SSIM. As an additional evaluation metric, we run the segmentation algorithm proposed by [79] on the ground truth and reconstruction image and compare the resulting segmentation masks using the SSIM. The resulting values are given in the following table.
@@ 7 Conclusions
This work can be extended in several directions. First, gradient descent methods converge only locally and are not necessarily fast. Therefore, it would be interesting to extend the minimization of the functional $F$ in Section 12 to higher-order methods or incorporate momentum parameters. Moreover, a careful choice of the initialization could improve the convergence behavior. Further, our reconstruction method could be extended to Bayesian inverse problems. Since the mixture model of VAEs provides us with a probability distribution and an (approximate) density, stochastic sampling methods like the Langevin dynamics could be used for quantifying uncertainties within our reconstructions. Indeed, Langevin dynamics on Riemannian manifolds are still an active area of research. Further, for large numbers $K$ of charts the mixture of VAEs might have a considerable number of parameters. As a remedy, we could incorporate the selection of the chart as conditioning parameter in one conditional decoder-encoder pair, see [90] as a reference for conditional VAEs. Finally, recent papers show that diffusion models provide an implicit representation of the data manifold ([14, 85]). It would be interesting to investigate optimization models on such manifolds in order to apply them to inverse problems.

### v3

@@ 4.2 Retractions for Learned Charts
In order to apply this gradient scheme for a learned manifold given by the learned mappings $(\mathcal{D}_{k},\mathcal{E}_{k})_{k=1}^{K}$, we consider two types of retractions. We introduce them and show that they are indeed retractions in the following lemmas. The first one generalizes the idea from Lemma 4 and Proposition 5 in Absil and Malick (2012) of moving along the tangent vector in $\mathbb{R}^{n}$ and reprojecting onto the manifold. However, the results by Absil and Malick (2012) are based on the orthogonal projection, which is hard or even impossible to compute. Thus, we replace it by some more general projection $\pi$. In our applications, $\pi$ will be chosen as in Remark 3, i.e., we set $\pi(x)=\mathcal{D}_{k}(\mathcal{E}_{k}(x)).$
Lemma 6
@@
Proof  The property $R_{x}(0)=x$ is directly clear from the definition of $R_{x}$. Now let $h\in T_{x}\mathcal{M}\subseteq\mathbb{R}^{n}$ and $\gamma_{h}\colon(-\epsilon,\epsilon)\to\mathcal{M}$ be a differentiable curve with $\gamma_{h}(0)=x$ and $\dot{\gamma}_{h}(0)=h$. As ${\pi}_{|U}$ is the identity on $\mathcal{M}$, we have by the chain rule that
@@
The second retraction uses the idea of changing to local coordinates, moving into the gradient direction by using the local coordinates and then going back to the manifold representation. Note that similar constructions are considered in Section 4.1.3 in the book of Absil et al. (2009). However, as we did not find an explicit proof for the lemma, we give it below for the sake of completeness.
Lemma 7
@@
Proof  The property $R_{x}(0)=x$ is directly clear from the definition of $R_{x}$. Now let $h\in T_{0}(T_{x}\mathcal{M})=T_{x}\mathcal{M}\subseteq\mathbb{R}^{n}$. By (17), we have that there exists some $y\in\mathbb{R}^{d}$ such that $h=Jy$. Then, we have by the chain rule that
@@ 4.3 Gradient Descent on Learned Manifolds
By Lemma 6 and 7, we obtain that the mappings
@@
Remark 8 (Descent algorithm)
By construction, Algorithm 3 is a descent algorithm. That is, for a sequence $(x_{n})_{n}$ generated by the algorithm it holds that $F(x_{n+1})\leq F(x_{n})$. With the additional assumption that $F$ is bounded from below, we have that $(F(x_{n}))_{n}$ is a bounded descending and hence convergent sequence. However, this does neither imply convergence of the iterates $(x_{n})_{n}$ themselves nor optimality of the limit of $(F(x_{n}))_{n}$. For more details on the convergence of line-search algorithms on manifolds we refer to Section 4 of the book by Absil et al. (2009).
@@ 5 Numerical Examples
Remark 9 (Dimension of the Manifold)
For all our numerical experiments, we assume that the dimension $d$ of the data manifold is known. This assumption might be violated for practical applications. However, there exist several methods in the literature to estimate the dimension of a manifold from data (see, e.g., Stanczuk et al., 2024; Camastra and Staiano, 2016; Fan et al., 2010; Levina and Bickel, 2004). We are aware that dimension estimation of high dimensional data sets is a hard problem which cannot be considered as completely solved so far. In particular, most of these algorithms make assumptions on the distribution of the given data points. Even though it is an active area of research, it is not the scope of this paper to test, benchmark or develop such algorithms. Similarly, combining them with our mixture of VAEs is left for future research.
@@ 6 Mixture of VAEs for Inverse Problems
In many inverse problems, the unknown $x$ can be modeled as an element of a low-dimensional manifold $\mathcal{M}$ in $\mathbb{R}^{n}$ (Alberti et al., to appear; Asim et al., 2020; Bora et al., 2017; Hyun et al., 2021; Ongie et al., 2020; Seo et al., 2019; Massa et al., 2022; Alberti et al., 2023), and this manifold can be represented by the mixture of VAEs as explained in Section 3. Thus, the solution of (20) can be found by optimizing the function
@@
Throughout these experiments we consider images of size $128\times 128$ and use the architecture from Section 3.2 with $L=3$. Starting with the latent dimension $d$, the mapping $A_{1}\colon\mathbb{R}^{d}\to\mathbb{R}^{32^{2}}$ fills up the input vector $x$ with zeros up to the size $32^{2}$, i.e., we set $A_{1}(x)=(x,0)$. The invertible neural network $T_{1}\colon\mathbb{R}^{32^{2}}\to\mathbb{R}^{32^{2}}$ consists of $3$ invertible blocks, where the subnetworks $s_{i}$ and $t_{i}$, $i=1,2$ are dense feed-forward networks with two hidden layers and $64$ neurons. Afterwards, we reorder the dimensions to obtain an image of size $32\times 32$. The mappings $A_{2}\colon\mathbb{R}^{32\times 32}\to\mathbb{R}^{32\times 32\times 4}$ and $A_{3}\colon\mathbb{R}^{64\times 64}\to\mathbb{R}^{64\times 64\times 4}$ copy each channel $4$ times. Then, the generalized inverses $A_{2}^{\dagger}\colon\mathbb{R}^{32\times 32\times 4}\to\mathbb{R}^{32\times 32}$ and $A_{3}^{\dagger}\colon\mathbb{R}^{64\times 64\times 4}\to\mathbb{R}^{64\times 64}$ from (13) are given by taking the mean of the four channels of the input. The invertible neural networks $T_{2}\colon\mathbb{R}^{32\times 32\times 4}\to\mathbb{R}^{64\times 64}$ and $T_{3}\colon\mathbb{R}^{64\times 64\times 4}\to\mathbb{R}^{128\times 128}$ consist of $3$ invertible blocks, where the subnetworks $s_{i}$ and $t_{i}$, $i=1,2$ are convolutional neural networks with one hidden layer and $64$ channels. After these three coupling blocks we use an invertible upsampling (Etmann et al., 2020) to obtain the correct output dimension. For the normalizing flow in the latent space, we use an invertible neural network with three blocks, where the subnetworks $s_{i}$ and $t_{i}$, $i=1,2$ are dense feed-forward networks with two hidden layers and $64$ neurons.
@@ 6.2 Electrical Impedance Tomography
Finally, we consider the highly non-linear and ill-posed inverse problem of electrical impedance tomography (EIT, Cheney et al., 1999), which is also known in the mathematical literature as the Calderón problem (Astala and Päivärinta, 2006; Feldman et al., 2019; Mueller and Siltanen, 2012). EIT is a non-invasive, radiation-free method to measure the conductivity of a tissue through electrodes placed on the surface of the body. More precisely, electrical currents patterns are imposed on some of these electrodes and the resulting voltage differences are measured on the remaining ones. Although harmless, the use of this modality in practice is very limited because the standard reconstruction methods provide images with very low spatial resolution. This is an immediate consequence of the severe ill-posedness of the inverse problem (Alessandrini, 1988; Mandache, 2001).
Classical methods for solving this inverse problem include variational-type methods (Cheney et al., 1990), the Lagrangian method (Chen and Zou, 1999), the factorization method (Brühl and Hanke, 2000; Kirsch and Grinberg, 2008), the D-bar method (Siltanen et al., 2000), the enclosure method (Ikehata and Siltanen, 2000), and the monotonicity method (Tamburrino and Rubinacci, 2002). Similarly as many other inverse problems, deep learning methods have had a big impact on EIT. For example, Fan and Ying (2020) propose an end-to-end neural network that learns the forward map $\mathcal{G}$ and its inverse. Moreover, deep learning approaches can be combined with classical methods, e.g, by post processing methods (Hamilton and Hauptmann, 2018; Hamilton et al., 2019) or by variational learning algorithms (Seo et al., 2019).
@@
Remark 10 (Dimension and topology of the data manifold)
@@
A slightly more general version of this manifold was considered by Alberti et al. (2023), where Lipschitz stability is proven for a related inverse boundary value problem restricted to the manifold. Other types of inclusions (with unknown locations), notably polygonal and polyhedral inclusions, have been considered in the literature (Beretta et al., 2021; Beretta and Francini, 2022; Aspri et al., 2022). The case of small inclusions is discussed by Ammari and Kang (2004).
@@
Since the PDE (22) is linear, the map $\mathcal{G}(\gamma)$ is linear. However, the forward map $\gamma\mapsto\mathcal{G}(\gamma)$ is nonlinear in $\gamma$, and so is the corresponding inverse problem. The map $\mathcal{G}$ is continuously differentiable, and its Fréchet derivative (see Harrach, 2019) is given by $[\nabla\mathcal{G}(\gamma)](\sigma)(g)=w_{g}\big{|}_{\partial\Omega}$, where $w_{g}\in H^{1}(\Omega)$ is the unique weak solution with zero boundary mean of
@@
In our implementations, we discretize the linear mappings $G(\gamma)$ by restricting them to a finite dimensional subspace spanned by a-priori fixed boundary functions $g_{1},\dots,g_{N}\in H^{-\frac{1}{2}}_{\diamond}(\partial\Omega)$. Then, following (Beretta et al., 2018, eqt. (2.2)), we reconstruct the conductivity by minimizing the semi-discrete functional
@@
where $(u_{\text{true}})_{g_{n}}$ is the observed data. In our discrete setting, we represent the conductivitiy $\gamma$ by a piecewise constant function $\gamma=\sum_{m=1}^{M}\gamma_{m}\mathbbm{1}_{T_{m}}$ on a triangulation $(T_{m})_{m=1,\dots,M}$. Then, following Equation (2.20) from Beretta et al. (2018), the derivative of (23) with respect to $\gamma$ is given by
@@
In our experiments the domain $\Omega$ is given by the unit square $[0,1]^{2}$. For solving the PDEs (22) and (25), we use a finite element solver from the DOLFIN library (Logg and Wells, 2010). We employ meshes that are coarser in the middle of $\Omega$ and finer close to the boundary. To simulate the approximation error of the meshes, and to avoid inverse crimes, we use a fine mesh to generate the observation and a coarser one for the reconstructions. We use $N=15$ boundary functions, which are chosen as follows. We divide each of the four edges of the unit square $[0,1]^{2}$ into 4 segments and denote by $b_{1},\dots,b_{16}$ the functions that are equal to $1$ on one of these segments and 00 otherwise. Then, we define the boundary functions as $g_{n}=\sum_{i=1}^{16}a_{n,i}b_{i}$, where the matrix $A=(a_{n,i})_{n=1,\dots,15,i=1,\dots,16}$ is the $16\times 16$ Haar matrix without the first row. More precisely, the rows of $A$ are given by the rows of the matrices $2^{-k/2}(\mathrm{Id}_{2^{4-k}}\otimes(1,-1)\otimes e_{2^{k-1}}^{\mathrm{T}})$ for $k=1,\dots,4$, where $\otimes$ is the Kronecker product and $e_{j}\in\mathbb{R}^{j}$ is the vector where all entries are $1$.
@@ Results.
The reconstructions for 20 different ground truths are visualized in Figure 9(d). We observe that both models capture the ground truth structure in most cases, but also fail sometimes. Nevertheless, the reconstructions with the mixture of two VAEs recover the correct structure more often and more accurately than the single VAE, which can be explained by the better coverage of the data manifold. To quantify the difference more in detail, we rerun the experiment with $200$ different ground truth images and compare the results with one and two generators using the PSNR and SSIM. As an additional evaluation metric, we run the segmentation algorithm proposed by Otsu (1979) on the ground truth and reconstruction image and compare the resulting segmentation masks using the SSIM. The resulting values are given in the following table.
@@ 7 Conclusions
This work can be extended in several directions. First, gradient descent methods converge only locally and are not necessarily fast. Therefore, it would be interesting to extend the minimization of the functional $F$ in Section 14 to higher-order methods or incorporate momentum parameters. Moreover, a careful choice of the initialization could improve the convergence behavior. Further, our reconstruction method could be extended to Bayesian inverse problems. Since the mixture model of VAEs provides us with a probability distribution and an (approximate) density, stochastic sampling methods like the Langevin dynamics could be used for quantifying uncertainties within our reconstructions. Indeed, Langevin dynamics on Riemannian manifolds are still an active area of research. Further, for large numbers $K$ of charts the mixture of VAEs might have a considerable number of parameters. As a remedy, we could incorporate the selection of the chart as conditioning parameter in one conditional decoder-encoder pair (see Sohn et al., 2015 as a reference for conditional VAEs). Finally, recent papers show that diffusion models provide an implicit representation of the data manifold (Stanczuk et al., 2024; Ross et al., 2024). It would be interesting to investigate optimization models on such manifolds in order to apply them to inverse problems.
Acknowledgments and Disclosure of Funding
This material is based upon work supported by the Air Force Office of Scientific Research under award numbers FA8655-20-1-7027 and FA8655-23-1-7083. We acknowledge the support of Fondazione Compagnia di San Paolo. Co-funded by the European Union (ERC, SAMPDE, 101041040). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. GSA, MS and SS are members of the “Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni”, of the “Istituto Nazionale di Alta Matematica”. The research of GSA, MS and SS was supported in part by the MUR Excellence Department Project awarded to the Department of Mathematics, University of Genoa, CUP D33C23001110001. JH acknowledges funding by the German Research Foundation (DFG) within the project STE 571/16-1 and by the EPSRC programme grant “The Mathematics of Deep Learning” with reference EP/V026259/1.